OpenStudy (anonymous):

The regular price of a child's entry ticket to a water park is $6 less than that for an adult's. The park offers half off all entry tickets during the off-peak season. The Jordans paid a total of $69 for 2 adult tickets and 1 child's ticket to the water park during the off-peak season. The following equation represents this situation, where x represents the regular price of an adult ticket. : 69 = x + one-half(x - 6) What is the regular price of a child's ticket?

OpenStudy (anonymous):

What is the regular price of a child's ticket? $38 $42 $48 $66

OpenStudy (anonymous):

$48 because: (one-half changed to decimal format) \[69=x+0.5(x-6)\]\[69=x+0.5x-3\]\[+3 +3\]\[72=1.5x\]\[\div1.5 \div1.5\]\[48=x\]

OpenStudy (anonymous):

Thank you so much!! @Ricky_Bobby

OpenStudy (anonymous):

you're welcome. do you need help with anything else?